OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)?

hero (hero):

\[m = \frac {y_2 -y_1} {x_2 - x_1}\]

hero (hero):

that's the formula for slope

hero (hero):

\[y-y_1 = m(x- x_1) \]

hero (hero):

that's the point slope formula

hero (hero):

What you do is, find the slope first, then enter one of the given points into the point slope formula along with m, the slope, then simplify